# On the volume of K-semistable Fano manifolds

### v1

@@ Abstract.
We prove that the volume of an $n$-dimensional K-semistable Fano manifold that is not ${\mathbb{P}}^{n}$ is at most $2n^{n}$. Moreover, the equality holds only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$ or $X$ is a smooth quadric hypersurface $Q\subset{\mathbb{P}}^{n+1}$. Our proof is based on a new connection between K-semistability and minimal rational curves. More generally, we show that the volume of a K-semistable Fano manifold with a minimal rational curve of degree $d$ is bounded above by the volume of ${\mathbb{P}}^{d-1}\times{\mathbb{P}}^{n-d+1}$.
@@ 1. Introduction
In this paper, we study the anticanonical volume (or degree) of K-semistable Fano varieties. For a Fano variety $X$, its (anticanonical) volume ${\rm vol}(X)$ is defined to be the self-intersection number $(-K_{X})^{n}$. It is known that the volume of an $n-$dimensional Fano manifold can exceed the volume of ${\mathbb{P}}^{n}$, even among toric Fano manifolds (see [Deb01, Section 5.11]). In a major advance in the study of K-stability, Fujita [Fuj18] proved that the volume of an $n$-dimensional K-semistable Fano manifolds $X$ satisfies ${\rm vol}(X)\leq(n+1)^{n}={\rm vol}({\mathbb{P}}^{n})$ and the first equality holds if and only if $X\cong{\mathbb{P}}^{n}$. It was later generalized to possibly singular ${\mathbb{Q}}$-Fano varieties by Liu [Liu18]. The toric Fano case was proved earlier in [BB17]. Moreover, from the boundedness of K-semistable ${\mathbb{Q}}$-Fano varieties ([Jia20, Corollary 1.2]), we know that the set of volumes of $n$-dimensional K-semistable ${\mathbb{Q}}$-Fano varieties is finite away from zero. In this paper, we solve the conjecture on characterizing the second-largest volume for K-semistable Fano manifolds,
@@
Any K-semistable Fano manifold $X$ that is not ${\mathbb{P}}^{n}$ satisfies ${\rm vol}(X)=(-K_{X})^{n}\leq 2n^{n}$ and the equality holds only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$ or $X$ is a smooth quadric hypersurface $Q\subset{\mathbb{P}}^{n+1}$. In particular, this holds for any Kähler-Einstein metric manifold with positive Ricci curvature.
Like the result of Fujita, the last statement could be seen as a result in Kähler differential geometry. However, our proof uses purely algebraic geometry. The above result seems to be first conjectured in [AIM20, Problem 2.6], and is stated as a folklore conjecture in [Zhu25, Conjecture 6.8] which highlights the interesting (but also mysterious) feature that there are two Fano manifolds with second largest volume. Our proof will indeed give a satisfactory explanation of this feature by connecting it to the theory of minimal rational curves on Fano manifolds.
There is a local analogue of this conjecture sometimes called the ODP conjecture (see Conjecture 2.3) for the volume of klt singularity (see [SS17, LX19]). The ODP conjecture would imply that the volume of singular K-semistable Fano varieties is strictly less than $2n^{n}$ (by Liu’s local-to-global volume comparison, see Remark 3.9). Conversely, using Theorem 1.1, we immediately verify the ODP conjecture for the Fano cone over a K-semistable smooth Fano manifold (see Theorem 4.5).
The toric version of Theorem 1.1 has its own interest and is called the “gap hypothesis” in [AB24, Conjecture 3.10]. It implies the sharp bound of the canonical height of canonical model of toric Fano varieties over ${\rm Spec}\,{\mathbb{Z}}$ by [AB24, Lemma 3.8]. Moreover, in the toric case, the ODP conjecture is solved by Moraga-Süß’s result ([MS24]) using convex geometric methods. So combined with Theorem 1.1, we indeed have the sharp upper bound for all possibly singular K-semistable toric Fano varieties.
@@
The second largest volume of $n$-dimensional K-semistable toric Fano varieties is $2n^{n}$ and the equality holds only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$.
Corollary 1.3.
Suppose $P\subseteq{\mathbb{R}}^{n}$ is a $n$-dimensional reflexive lattice polytope with barycenter at $0\in{\mathbb{R}}^{n}$. Assume $P$ is not unimodularly equivalent to $(n+1)$ times a standard simplex $(n+1)\Delta_{n}$, then the volume of $P$ with respect to the Lebesgue measure in ${\mathbb{R}}^{n}$ satisfies ${\rm vol}_{{\mathbb{R}}^{n}}(P)\leqslant 2n^{n}/n!$ and the equality holds if and only if $P$ is unimodularly equivalent to $[0,2]\times(n\Delta_{n-1})$.
We will prove Theorem 1.1 by exploring a new connection between K-stability and minimal rational curves. Roughly speaking, we are going to estimate the volume by using the valuative criterion to test K-semistability via valuations associated to weighted blowups along minimal rational curves. An interesting new phenomenon we find is that the volume estimates become sharp when we let certain weights go to infinity, and we indeed get the following general result, which to our knowledge is the first result connecting explicitly the theory of K-stability and the theory of minimal rational curves on Fano manifolds.
Theorem 1.4.
Assume that a K-semistable Fano manifold $X$ admits a minimal rational curve $f:{\mathbb{P}}^{1}\rightarrow X$ such that $f^{*}TX={\mathcal{O}}(2)\oplus{\mathcal{O}}(1)^{\oplus(d-2)}\oplus{\mathcal{O}}^{\oplus(n-d+1)}$ for $12\leq d\leq n+$ 1. Then we have ${\rm vol}(X)\leq{\rm vol}({\mathbb{P}}^{d-1}\times{\mathbb{P}}^{n-d+1})$.
The smooth toric case is special in the sense that there are always smoothly embedded projective spaces with trivial normal bundles (see [CFH14, Ara06]). So the toric Fano case can also be considered as an example of the following direct generalization of Fujita’s result.
Theorem 1.5.
Let $X$ be an $n$-dimensional K-semistable Fano manifold, and $Z\subset X$ a codimension $r$ smooth complex submanifold. Assume that the normal bundle of $Z$ inside $X$ is trivial and set $d=(-K_{X})^{n-r}\cdot Z$. Then we have the inequality
@@
Moreover, the equality holds if and only if $X$ is biholomorphic to ${\mathbb{P}}^{r}\times Z$. In particular $Z$ is also a K-semistable Fano manifold.
We sketch the organization of this paper. In the next section, we recall basic knowledge about some key concepts used in later sections: K-stability, Seshadri constant, and minimal rational curves. In section 3, we prove Theorem 1.5, and Theorem 1.1 in the case the Fano manifold contains a minimal rational curve with a trivial normal bundle. We also deduce Theorem 1.2 and Corollary 1.3. In section 4, we prove Theorem 1.4 to achieve the full proof of Theorem 1.1. We end the paper by giving some related examples and finding (K-semistable) Fano manifolds with minimal volumes.
@@ Acknowledgments
C. Li is partially supported by NSF (Grant No. DMS-2305296). M. Miao would like to thank his advisor Gang Tian for his constant support, encouragement, and guidance. M. Miao would like to thank Jiyuan Han, Xiaowei Jiang, Yang Liu, Yalong Shi, and Linsheng Wang for helpful discussions. We thank Robert Berman, Kento Fujita, Yuchen Liu, Chenyang Xu and Ziquan Zhuang for their interest and helpful comments, and Kento Fujita for pointing to us the Fact 1. This work is carried out during M. Miao’s visit to Rutgers University. He would like to thank the graduate school of Nanjing University for providing funding support and thank the Department of Mathematics at Rutgers University for its hospitality.
@@ 2. Preliminaries
We will work over ${\mathbb{C}}$. Unless otherwise specified, all varieties are assumed to be normal and projective. A ${\mathbb{Q}}$-Fano variety $X$ is a variety with at worst klt singularities such that the anti-canonical divisor $-K_{X}$ is an ample ${\mathbb{Q}}$-Cartier divisor. A singularity $x\in X$ consists of a variety $X$ and a closed point $x\in X$. A singularity $x\in X$ is called klt if $X$ is klt near a neighborhood of $x$. A ${\mathbb{R}}$-valuation over singularity $x$ is a valuation $v\colon K(X)^{*}\rightarrow{\mathbb{R}}$ center at $x$ (namely, for all $f\in{\mathcal{O}}_{X,x}$, we have $v(f)\geqslant 0$ and $v(f)>0$ if and only if $f\in\mathfrak{m}_{x}$) and $v|_{{\mathbb{C}}^{*}}=0$. The set of all such valuations is denoted by ${\rm Val}_{X,x}$.
@@ 2.1. K-stability
K-stability, first introduced by Tian ([Tia97]) and later reformulated algebraically by Donaldson ([Don02]), is an algebro-geometric notion to characterize the existence of Kähler-Einstein metrics on Fano varieties. In this subsection, we recall some notions in K-stability theory that are relevant to our paper, and refer to [Xu24] for a detailed exposition of K-stability theory.
@@
$S(E):=S(-K_{X};E)\coloneqq\frac{1}{(-K_{X})^{n}}\int_{0}^{\infty}{\rm vol}(-K_{X})-xE)\,{\rm d}x,$
@@
Another equivalent way to characterize K-semistability is via the $\delta$-invariant (also known as stability threshold), we recall $\delta(X)\coloneqq{\rm inf}_{E/X}\frac{A(E)}{S(E)}$ where the infinum is taking over all divisors $E$ over $X$ ([FO18, BJ20]). Then $X$ is K-semistable if and only if $\delta(X)\geqslant 1$ by the valuative criterion.
@@
(ODP conjecture) The second largest local volume of an $n$-dimensional klt singularity is $2(n-1)^{n}$, with equality if and only if $x\in X$ is an ordinary double point.
@@
(2) If $\psi\in\mathscr{F}$ satisfies $(\phi^{-1})^{\prime}(V)\geq(\psi^{-1})^{\prime}(V)$ for any $V\in(0,\infty)$, then $F^{\phi}(V)\geq F^{\psi}(V)$ for any $V\in(0,\infty)$. If moreover the first strict inequality holds for $V\geq V_{1}$, then $F^{\phi}(V)>F^{\psi}(V)$ for $V\geq V_{1}$. As a consequence, we get $(F^{\phi})^{-1}(A)<(F^{\psi})^{-1}(A)$ for any $A\in(0,\infty)$ with strict inequality if $A>F^{\psi}(V_{1})$.
@@
First assume that $\phi(x)$ is smooth and strictly increasing on $[0,\infty)$. Then its inverse function $\phi^{-1}(x)$ is also smooth, strictly increasing and satisfies (2). Set $G(V)=\int_{0}^{\phi^{-1}(V)}(V-\phi(x)){\rm d}x$ so that $F(V)=F^{\phi}(V)=G(V)/V$. Then $G$ is a smooth function satisfying $G(0)=0$. Its derivative is equal to $G^{\prime}(V)=\int_{0}^{\phi^{-1}(V)}1\,{\rm d}x=\phi^{-1}(V)$. We can then calculate the derivative of $F(V)$ at any $V\in(0,\infty)$:
@@
If $\phi$ is only piecewise smooth, we can just carry out the above argument piecewise on each interval of smoothness. The same argument applies to the second statement too. ∎
@@
There exists a unique solution $T=T(A_{X}(E))$ to the equation
@@
By the above lemma, $F^{\phi}(V)$ is a strictly increasing function that diverges to +$\infty$+ ∞ as $V\rightarrow+\infty$. So there exists a unique $V^{*}$ that satisfies $A(E)=F^{\phi}(V^{*})$ which is equivalent to the equality:
@@
Assume that $\phi(x)=2nx^{n-1}$ and $A(E)=n-1$. The $\Phi(x)=2x^{n}$. The equation (4) has the solution $T=n$ and $\phi(T)=2n^{n}$.
@@ 2.2. Seshadri Constant
is an ample divisor since $(m\epsilon+r)/(m+1)<\epsilon$. Then, by Kodaira’s vanishing theorem, we get $H^{1}(\hat{X},mB-E)=0$. Therefore, the natural map $H^{0}(\hat{X},mB)\rightarrow H^{0}(E,mB|_{E})$ is surjective for all $m>0$ such that $mB$ is Cartier. We conclude that $\Phi|_{E}\colon E\rightarrow Y$ is a closed embedding. Since $\hat{X}$ has klt singularites, then it has rational singularities ([KM98, Theorem 5.22]), so in particular it is Cohen-Macaulay. Since $B=-K_{\hat{X}}+(r-1-\epsilon)E\sim_{\Phi,{\mathbb{Q}}}0$, we get $-K_{\hat{X}}\sim_{\Phi,{\mathbb{Q}}}\lambda E$ where $\lambda-r+1+\epsilon>1$. By [LZ18, Lemma 8], we have $\Phi\colon\hat{X}\rightarrow Y$ is not birational. Thus, $\Phi$ must be a fiber type contraction.
@@
$0\rightarrow{\mathcal{O}}_{Y}\rightarrow{\mathcal{E}}\rightarrow{\mathcal{O}}_{Y}(-1)\rightarrow 0$

### v2

@@ Abstract.
We prove that the anti-canonical volume of an $n$-dimensional K-semistable Fano manifold that is not ${\mathbb{P}}^{n}$ is at most $2n^{n}$. Moreover, the volume is equal to $2n^{n}$ if and only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$ or $X$ is a smooth quadric hypersurface $Q\subset{\mathbb{P}}^{n+1}$. Our proof is based on a new connection between K-semistability and minimal rational curves.
@@ 1. Introduction
In this paper, we study the anticanonical volume (or degree) of K-semistable Fano varieties. For a Fano variety $X$, its (anticanonical) volume ${\rm vol}(X)$ is defined to be the self-intersection number $(-K_{X})^{n}$. It is known that the volume of an $n$-dimensional Fano manifold can exceed the volume of the complex projective space ${\mathbb{P}}^{n}$, even among toric Fano manifolds (see [Deb01, Section 5.11]). In a major advance in the study of K-stability, Fujita [Fuj18] proved that the volume of an $n$-dimensional K-semistable Fano manifold $X$ satisfies ${\rm vol}(X)\leq(n+1)^{n}={\rm vol}({\mathbb{P}}^{n})$ and the first equality holds if and only if $X\cong{\mathbb{P}}^{n}$. It was later generalized to possibly singular $\mathbb{Q}$}-Fano varieties by Liu [Liu18]. The toric Fano case was proved earlier in [BB17]. Moreover, from the boundedness of K-semistable $\mathbb{Q}$}-Fano varieties ([Jia20, Corollary 1.2]), we know that the set of volumes of $n$-dimensional K-semistable $\mathbb{Q}$}-Fano varieties is finite away from zero. In this paper, we solve the conjecture on characterizing the second-largest volume for K-semistable Fano manifolds,
@@
Any K-semistable Fano manifold $X$ that is not ${\mathbb{P}}^{n}$ satisfies $(-K_{X})^{n}\leq 2n^{n}$ and the equality holds only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$ or $X$ is a smooth quadric hypersurface $Q\subset{\mathbb{P}}^{n+1}$. In particular, this holds for any Kähler manifold admitting Kähler-Einstein metric with positive Ricci curvature.
Like the result of Fujita ([Fuj18]), the last statement could be seen as a result in Kähler geometry. However, our proof uses purely algebraic geometry. The above result seems to be first conjectured in [AIM20, Problem 2.6], and is stated as a folklore conjecture in [Zhu25, Conjecture 6.8] which highlights the interesting (but also mysterious) feature that there are two Fano manifolds with the second largest volume. Our proof will indeed give a satisfactory explanation of this feature by connecting it to the theory of minimal rational curves on Fano manifolds.
There is a local analogue of this conjecture sometimes called the ODP conjecture (see Conjecture 2.3) for the volume of klt singularity (see [SS17, LX19]). The ODP conjecture would imply that the volume of singular K-semistable Fano varieties is strictly less than $2n^{n}$ (by Fujita-Liu’s local-to-global volume comparison, see Remark 3.7). Conversely, using Theorem 1.1, we immediately verify the ODP conjecture for the Fano cone over a K-semistable smooth Fano manifold (see Theorem 4.15).
The toric version of Theorem 1.1 has its own interest and is called the “gap hypothesis” in [AB24, Conjecture 3.10]. It implies the sharp bound of a canonical height of canonical model of toric Fano varieties over $\rm Spec$$\mathbb{Z}$} by [AB24, Lemma 3.8]. Moreover, in the toric case, the ODP conjecture is solved by Moraga-Süß’s result ([MS24]) using convex geometric methods. So combined with Theorem 1.1, we indeed have the sharp upper bound for all possibly singular K-semistable toric Fano varieties.
Theorem 1.2 (See Theorem 3.6).
Let $X$ be an $n$-dimensional K-semistable toric Fano variety that is not ${\mathbb{P}}^{n}$. Then $(-K_{X})^{n}\leq 2n^{n}$ and the equality holds only if $X\cong{\mathbb{P}}^{1}\times{\mathbb{P}}^{n-1}$.
Corollary 1.3 (See Corollary 3.9).
Suppose $P\subseteq{\mathbb{R}}^{n}$ is an $n$-dimensional reflexive lattice polytope with its barycenter at $0\in{\mathbb{R}}^{n}$. Assume $P$ is not unimodularly equivalent to $(n+1)$ times a standard simplex $(n+1)\Delta_{n}$. Then the volume of $P$ with respect to the Lebesgue measure in ${\mathbb{R}}^{n}$ satisfies ${\rm vol}_{{\mathbb{R}}^{n}}(P)\leqslant 2n^{n}/n!$ and the equality holds if and only if $P$ is unimodularly equivalent to $0,2]\times(n\Delta_{n-1})$.
We will prove Theorem 1.1 by exploring a new connection between K-stability and minimal rational curves. We use the following invariant
$l_{X}={\rm min}\{(-K_{X}\cdot C)|\,C\subset X\,\text{is a minimal rational curve on}\,X\}\in\{2,3,\cdots,\dim X+1\}.$
For the case of $l_{X}=\dim X+1$ and $l_{X}=\dim X$, we can use the classification result of [CMSB02, Miy04, CD15, DH17] to verify Theorem 1.1. At the other extreme, if $l_{X}=2$ and when the minimal rational curve is embedded, then the minimal rational curve has a trivial normal bundle. In this special case Theorem 1.1 is a consequence of the following generalization of a result of K. Fujita ([Fuj18]).
@@
Let $X$ be an $n$-dimensional K-semistable Fano manifold, and $Z\subset X$ be a codimension $r$ smooth complex submanifold. Assume that the normal bundle of $Z$ inside $X$ is trivial and set $d=(-K_{X})^{n-r}\cdot Z$. Then we have the inequality
@@
Moreover, the equality holds if and only if $X$ is biholomorphic to ${\mathbb{P}}^{r}\times Z$. In particular in the equality case $Z$ is also a K-semistable Fano manifold.
To estimate the volume in much more involved cases when $3\leq d=l_{X}\leq n-1$, we are going to use the valuative criterion to test K-semistability via special valuations centered at minimal rational curves. While the proof of Theorem 1.4 uses ordinary blowup of submanifolds, the proof of Theorem 1.1 needs weighted blowups (see Remark 4.11). In particular, we will define and analyze a particular weighted blowup of type $({1}^{\oplus(d-2)},2^{\oplus(n-d+1)})$ that is adapted to the splitting of normal bundle over ${\mathbb{P}}^{1}$ (see (12)). The analysis will also allow us to overcome the difficulty that the minimal rational curve may have singularities.
We sketch the organization of this paper. In the next section, we recall basic knowledge about some key concepts used in later sections: K-stability, Seshadri constant, and minimal rational curves. In section 3, we prove Theorem 1.4 which implies Theorem 1.1 in the special case when the Fano manifold contains an embedded minimal rational curve with a trivial normal bundle. We also deduce Theorem 1.2 and Corollary 1.3. In section 4, we give the full proof of Theorem 1.1. We end the paper by giving some related examples and finding (K-semistable) Fano manifolds with minimal volumes.
@@ Acknowledgments
C. Li is partially supported by NSF (Grant No. DMS-2305296). M. Miao would like to thank his advisor Gang Tian for his constant support, encouragement, and guidance. M. Miao would like to thank Jiyuan Han, Xiaowei Jiang, Yang Liu, Yalong Shi, and Linsheng Wang for helpful discussions. We thank Robert Berman, Kento Fujita, Yuchen Liu, Chenyang Xu and Ziquan Zhuang for their interest and helpful comments, and Kento Fujita for pointing to us the Fact 1. We are especially grateful to Ziquan Zhuang for kindly pointing out some error in our estimates that appeared in the first version of this preprint (see Remark 4.4) and very helpful suggestions. This work is carried out during M. Miao’s visit to Rutgers University. He would like to thank the graduate school of Nanjing University for providing funding support and thank the Department of Mathematics at Rutgers University for its hospitality.
@@ 2. Preliminaries
We will work over $\mathbb{C}$}. Unless otherwise specified, all varieties are assumed to be normal and projective. A $\mathbb{Q}$}-Fano variety $X$ is a variety with at worst klt singularities such that the anti-canonical divisor $K_{X}$ is an ample $\mathbb{Q}$}-Cartier divisor. A singularity $x\in X$ consists of a variety $X$ and a closed point $x\in X$. A singularity $x\in X$ is called klt if $X$ is klt near a neighborhood of $x$. A $\mathbb{R}$}-valuation over singularity $x$ is a valuation $v\colon K(X)^{*}\rightarrow{\mathbb{R}}$ centered at $x$ (namely, for all $f\in{\mathcal{O}}_{X,x}$, we have $v(f)\geqslant 0$ and $v(f)>0$ if and only if $f\in\mathfrak{m}_{x}$) and $v|_{{\mathbb{C}}^{*}}=0$. The set of all such valuations is denoted by ${\rm Val}_{X,x}$.
@@ 2.1. K-stability
K-stability, first introduced by Tian ([Tia97]) and later reformulated algebraically by Donaldson ([Don02]), is an algebro-geometric notion to characterize the existence of Kähler-Einstein metrics on Fano varieties. In this subsection, we recall some notions in K-stability theory that are relevant to our paper, and refer to [Xu25] for a detailed exposition of K-stability theory.
@@
$S(E):=S(-K_{X};E)\coloneqq\frac{1}{(-K_{X})^{n}}\int_{0}^{\infty}{\rm vol}(-K_{X}-xE)\,{\rm d}x,$
@@
Another equivalent way to characterize K-semistability is via the $\delta$-invariant (also known as the stability threshold), we recall $\delta(X)\coloneqq{\rm inf}_{E/X}\frac{A(E)}{S(E)}$ where the infinum is taking over all divisors $E$ over $X$ ([FO18, BJ20]). Then $X$ is K-semistable if and only if $\delta(X)\geqslant 1$ by the valuative criterion.
@@
(ODP conjecture) The second largest local volume of an $n$-dimensional klt singularity is $2(n-1)^{n}$. Moreover, the local volume is equal to $2(n-1)^{n}$ if and only if $x\in X$ is an ordinary double point.
@@
(2) If $\psi\in\mathscr{F}$ satisfies $(\phi^{-1})^{\prime}(V)\geq(\psi^{-1})^{\prime}(V)$ for any $V\in(0,\infty)$, then $F^{\phi}(V)\geq F^{\psi}(V)$ for any $V\in(0,\infty)$. If moreover the first strict inequality holds for $V\geq V_{1}$, then $F^{\phi}(V)>F^{\psi}(V)$ for $V\geq V_{1}$. As a consequence, we get $(F^{\phi})^{-1}(A)\leq(F^{\psi})^{-1}(A)$ for any $A\in(0,\infty)$ with a strict inequality if $A>F^{\psi}(V_{1})$.
@@
First assume that $\phi(x)$ is smooth and strictly increasing on $0,\infty)$. Then its inverse function $\phi^{-1}(x)$ is also smooth, strictly increasing and satisfies (2). Set $G(V)=\int_{0}^{\phi^{-1}(V)}(V-\phi(x)){\rm d}x$ so that $F(V)=F^{\phi}(V)=G(V)/V$. Then $G$ is a smooth function satisfying $G(0)=0$. Its derivative is equal to $G^{\prime}(V)=\int_{0}^{\phi^{-1}(V)}1\,{\rm d}x=\phi^{-1}(V)$. We can then calculate the derivative of $F(V)$ at any $V\in(0,\infty)$:
@@
If $\phi$ is only piecewise smooth, we can just carry out the above argument piecewise on each interval of smoothness. The same argument applies to prove the second statement too. ∎
@@
Then there exists a unique solution $T=T(A_{X}(E))$ to the equation
@@
By the above lemma, $F^{\phi}(V)$ is a strictly increasing function that diverges to $\infty$ as $V\rightarrow+\infty$. So there exists a unique $V^{*}$ that satisfies $A(E)=F^{\phi}(V^{*})$ which is equivalent to the equality:
@@
Assume that $\phi(x)=2nx^{n-1}$ and A(E)=n1A(E)=n-1. The $\Phi(x)=2x^{n}$. The equation (4) has the solution $T=n$ and $\phi(T)=2n^{n}$.
@@ 2.2. Seshadri Constant
is an ample divisor since $(m\epsilon+r)/(m+1)<\epsilon$. Then, by Kodaira’s vanishing theorem, we get $H^{1}(\hat{X},mB-E)=0$. Therefore, the natural map $H^{0}(\hat{X},mB)\rightarrow H^{0}(E,mB|_{E})$ is surjective for all $m>0$ such that $mB$ is Cartier. We conclude that $\Phi|_{E}\colon E\rightarrow Y$ is a closed embedding. Since $\hat{X}$ has klt singularites, then it has rational singularities ([KM98, Theorem 5.22]), so in particular it is Cohen-Macaulay. Since $B=-K_{\hat{X}}+(r-1-\epsilon)E\sim_{\Phi,{\mathbb{Q}}}0$, we get $K_{\hat{X}}\sim_{\Phi,{\mathbb{Q}}}\lambda E$ where $\lambda-r+1+\epsilon>1$. By [LZ18, Lemma 8], we have $\Phi\colon\hat{X}\rightarrow Y$ is not birational. Thus, $\Phi$ must be a fiber type contraction.
@@
$0\rightarrow{\mathcal{O}}_{Y}\rightarrow{\mathcal{E}}\rightarrow{\mathcal{O}}_{Y}(-1)\rightarrow 0.$